# Spin Seebeck Effect as a Probe for Majorana Fermions in Kitaev Spin Liquids

### v1

@@ Abstract
Quantum entanglement in strongly correlated electron systems often leads to exotic elementary excitations. Quantum spin liquids provide a paradigmatic example, where the elementary excitations are described by fractional quasiparticles such as spinons. However, such fractional quasiparticles behave differently from electrons, making their experimental identification challenging. Here, we theoretically investigate the spin Seebeck effect, which is a thermoelectric response via a spin current, as an efficient probe of the fractional quasiparticles in quantum spin liquids, focusing on the Kitaev honeycomb model. By comprehensive studies using the real-time dynamics, the perturbation theory, and the linear spin-wave theory based on the tunnel spin-current theory, we find that the spin current is induced by thermal gradient in the Kitaev spin liquid, via the low-energy fractional Majorana excitations. This underscores the ability of Majorana fermions to carry spin current, despite lacking spin angular momentum. Furthermore, we find that the induced spin current changes its sign depending on the sign of the Kitaev interaction, indicating that the Majorana fermions contribute to the spin current with (up-)down-spin like nature when the exchange coupling is (anti)ferromagnetic. Thus, in contrast to the negative spin current already found in a one-dimensional quantum spin liquid, our finding reveals that the spin Seebeck effect can exhibit either positive or negative signals, contingent upon the nature of fractional excitations in the quantum spin liquids. We also clarify contrasting field-angle dependence between the Kitaev spin liquid in the low-field limit and the high-field ferromagnetic state, which is useful for the experimental identification. Our finding suggests that the spin Seebeck effect could be used not only to detect fractional quasiparticles emerging in quantum spin liquids but also to generate and control them.
@@ I Introduction
Quantum entanglement is the key to the emergence of topological phases which have been of central interest in strongly correlated electron systems. The topological phases can be regarded as vacuums of fractionalized elementary excitations which behave differently from the original electrons. A representative example is the fractional quantum Hall states in two-dimensional electron systems under a magnetic field [1], where electrons and magnetic fluxes form composite elementary excitations with fractional charges obeying neither the Bose-Einstein nor Fermi-Dircac statistics, called non-Abelian anyons [2, 3, 4]. Numerous studies have been conducted to identify and manipulate the non-Abelian anyons [5], since they are expected to be useful for future quantum technologies, such as fault-tolerant quantum computing [6, 7].
@@
The Kitaev spin liquid (KSL) is a rare quantum spin liquid that has been rigorously shown to be the ground state of a realistic quantum spin model on a two-dimensional honeycomb lattice [24]. The exact solution allows one to scrutinize the fundamental properties of fractional excitations. In the KSL, spins are fractionalized into itinerant Majorana fermions and localized $Z_{2}$ fluxes at zero magnetic field, and these two form a composite obeying anyonic statistics under a magnetic field. Stimulated by the proposal for realization of the model [25], tremendous efforts have been devoted to exploring candidate materials [26, 27, 28]. In addition, close collaborations between experiment and theory have unveiled exotic properties of the KSL [29, 30, 26, 31, 32], for instance, the half-quantized thermal Hall effect as direct evidence of the fractional Majorana excitations [33, 34, 35, 36, 37]. Recently, there are several proposals for generating, detecting, and controlling the fractional excitations in the KSL, e.g., by using the scanning tunneling microscopy [38, 39, 40, 41, 42], interferometers [43, 44, 45], impurity effects [46, 47], local lattice distortions [48], and magnetic fields [49], whose experimental demonstrations are awaited.
@@
Figure 1: (a) Schematic of an experimental setup for measuring the spin Seebeck effect in Kitaev magnets. A Kitaev magnet placed on a substrate is attached to heatbaths from two sides to generate a temperature gradient. A metal with strong spin-orbit coupling, e.g., Pt, is sandwiched at one end. Due to the inverse spin Hall effect, the spin current is converted into a voltage perpendicular to the thermal gradient in the metal, which is observable. (b) Kitaev model on a honeycomb lattice used in simulations. The black dots represent the spin-1/2 degrees of freedom, and the red, blue, and green lines represent the $x$, $y$, and $z$ bonds, respectively [see Eq. (1)]. The lattice with $X=4$ and $L=12$ is shown, where $X$ is the number of spins in the vertical direction and $L$ is those in the horizontal direction. We impose an open boundary condition in the horizontal direction and a periodic boundary condition in the vertical direction. The $a$, $b$, and $c$ axes and $S^{x}$, $S^{y}$, and $S^{z}$ axes are shown in the inset. $j_{0}$ denotes the site where the local dynamical spin susceptibility is calculated.
@@
The spin Seebeck effect refers to a phenomenon in which a spin current is induced by a thermal gradient. This effect is widely used as a sensitive probe for elementary excitations in magnets [50, 51, 52, 53, 54, 55, 56, 57]. In fact, in ordinary ferromagnets and antiferromagnets, the spin Seebeck effect is induced by magnon excitations from the conventional magnetic ordered state. Recently, this probe has been extended to more nontrivial magnets, such as those with frustrated interactions and low dimensionality, whose elementary excitations may have both positive and negative spin angular momenta, or even become spinless. The representative example is a negative spin Seebeck effect observed in a quantum spin liquid in a quasi-one-dimensional antiferromagnet [58]. This peculiar behavior, in contrast to conventional ferromagnets, was ascribed to fractional spinon excitations forming the Tomonaga–Luttinger liquid. Besides, other systems with strong quantum fluctuations, such as a spin-nematic liquid [59], a spin-Peierls system [60], and a magnon Bose–Einstein condensation system [61], have also been reported to exhibit the peculiar spin Seebeck effect. To the best of our knowledge, however, no study of quantum spin liquids in more than one dimension has been reported thus far. Recently, unconventional spin transport mediated by Majorana excitations in the KSL was reported [62, 63, 64, 65]. Meanwhile, as mentioned above, Majorana excitations were detected by thermal Hall measurements [33, 34, 35, 36, 37]. However, it remains unclear how the spin Seebeck effect, which involve both spin and heat transport, is manifested in the KSL and whether it is useful to identify, generate, and control the fractional excitations.
@@
In this paper, we theoretically analyze the spin current induced by a thermal gradient via the spin Seebeck effect in the KSL. The possible experimental setup is shown in Fig. 1(a); we consider a bilayer junction of a Kitaev magnet and a paramagnetic metal, and calculate the tunnel spin current through the interface by the local dynamical spin susceptibility in an applied magnetic field, based on the tunnel spin-current theory [66, 67]. Accurately calculating the dynamics of quantum many-body problems, particularly in more than one spatial dimension, poses a significant challenge. In addressing this issue, we opt for a state-of-the-art numerical technique, real-time dynamics simulations based on the time-dependent variational principle (TDVP) [68] with matrix product states (MPSs) [69, 70]. We also complementarily conduct the analysis based on the perturbation theory with respect to the magnetic field for the KSL, as well as the linear spin-wave theory for the field-induced ferromagnetic (FFM) state. As a result, we find that the spin current is induced by a thermal gradient in the KSL mediated by the low-energy fractional Majorana quasiparticles that have no spin angular momentum. This is qualitatively different from the fact that fractional spinons in forming a one-dimensional Tomonaga–Luttinger liquid induce a negative spin Seebeck effect since the spinon has spin-$1/2$. It is remarkable that the quasiparticles without angular momentum can contribute to the tunnel spin-current transport. Furthermore, we find that the sign of the tunnel spin current depends on the sign of the Kitaev interaction; namely, the ferromagnetic (FM) and AFM KSLs lead to a positive and negative spin Seebeck effect, respectively. This is in stark contrast to the behavior in conventional ferromagnetic states where the magnon excitations lead to a positive spin Seebeck effect. In addition, we find that the KSLs show distinctive field angle dependence of the spin current that is different from those of the FFM states. These peculiar behaviors, which are ascribed to the spin current mediated by fractional Majorana excitations, are revealed for the first time to our knowledge by using the state-of-the-art numerical technique. In turn, our findings suggest the possibility of generating and controlling the Majorana excitations by spin injection.
@@
The structure of this paper is as follows. In Sec. II, we describe the model and method used in this study. After introducing the Kitaev model under an external magnetic field, we describe the tunnel spin-current theory and numerical methods of real-time dynamics simulations to compute the spin Seebeck effect. We also mention the choice of model parameters and the details of calculation conditions. In Sec. III, we present the results of the real-time dynamics simulations, the perturbation analysis, and the linear spin-wave theory. In Sec. IV, we discuss the contrasting behaviors of the tunnel spin current depending on the sign of the Kitaev interaction, and the strength and direction of the magnetic field, which are important for experimental confirmation. Finally, Sec. V is devoted to summary.
@@ II Model and method
We investigate the spin Seebeck effect in the Kitaev model in Eq. (1) based on the tunnel spin-current theory [66, 67, 60, 71]. The experimental setup is shown in Fig. 1(a); the spin Seebeck effect is measured by the inverse spin Hall effect [72, 73, 74] in a metal with strong spin-orbit coupling, such as Pt, attached to a Kitaev magnet. In the tunnel spin-current theory, spin injection is assumed to occur locally through the interface between the metal and the magnet, and calculated under the following conditions: (i) The interaction between the metal and the magnet is small enough to be dealt with as a perturbation and conserves angular momenta, (ii) the polarization of the tunnel spin current is parallel to the direction of the magnetic field, (iii) the temperature difference between the metal and the magnet is small, and (iv) the metal is nonmagnetic with featureless density of states in the low-energy region. It is worth noting that although the total magnetization is not conserved in the Kitaev model, these assumptions allow us to evaluate the tunnel spin current transferred to the metal that is detected in experiments.
On these assumptions, the tunnel spin current induced by a thermal gradient is evaluated by the perturbation theory with respect to the interface interaction using the nonequilibrium Green function method [75, 76, 77]. The expression is given by
@@
up to a constant coefficient proportional to the temperature difference between the metal and magnet. The local dynamical spin susceptibility of the magnet, $\chi^{-+}_{\rm loc}(\omega)$, is defined as
@@
with the dumping factor $\delta$; $S^{\pm}_{j_{0}}$ is defined according to the direction of the magnetic field, namely,
@@
where $\tilde{x}$ and $\tilde{y}$ represent the directions perpendicular to the magnetic field along $\tilde{z}$ (the $\tilde{x}\tilde{y}\tilde{z}$ coordinate form a right-handed system) 111 We define $j_{0}$ on only one of the two sublattices as shown in Fig. 1(b) and neglect the sublattice dependence because trial simulation results do not show recognizable difference in $\chi_{\rm loc}^{-+}(\omega)$ when $j_{0}$ is defined on the other sublattice. . In Eq. (2), the integral kernel $k(x=\omega/T)$ is given by
@@
In this study, we compute the tunnel spin current in an applied magnetic field ${\bf B}$ along the $\tilde{z}(=z$, $a$, $b$, or $c)$ axis [see Fig. 1(b)], and mainly discuss the cases of ${\bf B}\parallel z$ and ${\bf B}\parallel c$; the former gives the Zeeman coupling as $-BS_{j}^{z}$, while the latter gives $-B(S_{j}^{x}+S_{j}^{y}+S_{j}^{z})/\sqrt{3}$ in the second term in Eq. (1). In the calculations of $X_{+}(\omega)$, we choose $S_{j_{0}}^{\pm}$ as $S_{j_{0}}^{\pm}=S_{j_{0}}^{x}\pm iS_{j_{0}}^{y}$ for ${\bf B}\parallel z$, and $S_{j_{0}}^{\pm}=S_{j_{0}}^{a}\pm iS_{j_{0}}^{b}$ for ${\bf B}\parallel c$ (cyclic permutations for ${\bf B}\parallel a$ and ${\bf B}\parallel b$). We investigate the low-field KSL as well as the high-field FFM state for comparison. The phase transition from the KSL to the FFM state was studied by using the exact diagonalization, the density matrix renormalization group (DMRG) [83, 84], and Majorana mean-field approximation [79, 80, 81, 82]. The phase diagram is different between the FM and AFM cases not only quantitatively but also qualitatively, as shown in Fig. 3. In the FM case, the KSL turns into the FFM state at $B\approx 0.03$ for both ${\bf B}\parallel c$ [79, 80, 82] and ${\bf B}\parallel z$ [81]; there is a single phase transition. Meanwhile, in the AFM case, the values of the critical field become one order of magnitude larger, and moreover, the system exhibits another phase transition in addition to that to the FFM state. For ${\bf B}\parallel c$, the KSL is stable up to $B\approx 0.4$, and a different spin liquid (SL) takes place before entering into the FFM state at $B\approx 0.6$ [80, 82]. For ${\bf B}\parallel z$, the KSL is stable up to $B\approx 0.4$, and the intermediate SL is predicted to be stable up to $B\approx 0.5$ [81]. With reference to these studies, we choose the values of $B$ for the KSL (FMM) state, $B_{\rm KSL}$ ($B_{\rm FFM}$), well below (above) the critical fields: $B_{\rm KSL}=0.01$ and $B_{\rm FFM}=0.5$ for the FM case, and $B_{\rm KSL}=0.1$ and $B_{\rm FFM}=1.5$ for the AFM case, as indicated by the red arrows in Fig. 3.
@@
In the following calculations, we compute $\chi^{-+}_{\rm loc}(\omega)$ in Eq. (3) at zero temperature for simplicity. This leaves the temperature dependence of the tunnel spin current only in the kernel $k(\omega/T)$. This assumption should be justified at sufficiently low temperature, especially below the crossover temperature $T\approx 0.016$ where the $Z_{2}$ flux excitations are almost absent [85]. To compute $\chi^{-+}_{\rm loc}(\omega)$ at $T=0$, we perform a real-time dynamics simulations based on TDVP [86]. The initial states of the simulations are obtained by multiplying $S^{\pm}_{j_{0}}$ to the MPS ground states computed by DMRG with the bond dimension $400$. We consider a cylinder-shaped lattice structure with the periodic (open) boundary condition in the vertical (horizontal) direction, as shown in Fig. 1(b), of the system size $X\times L$ with $(X,L)=(4,12)$, $(4,16)$, $(6,6)$, $(6,8)$, and $(6,10)$, where $X$ and $L$ are the number of sites in the vertical and horizontal directions, respectively, and employ the MPS representation wrapping around the cylinder in a snake form. We note that the ground state is not degenerate for all cases at the chosen magnetic fields. In Eq. (3), we choose $j_{0}$ at the site close to the center of each cylinder, as shown in Fig. 1(b), and take $\delta=0.01$. We set one time step of real-time dynamics as $\Delta t=0.2$ and perform $N_{t}=10^{4}$ time evolutions: Simulations are performed up to a maximum time $t_{\rm max}=2000$. When performing the Fourier transform [Eq. (3)], we interpolate the data by a cubic spline. The calculations are performed using a software library iTensor [87]. In addition, to understand the simulation results, we analyze the dynamical spin susceptibility based on the perturbation theory with respect to the magnetic field and the linear spin-wave theory, as described in Secs. III.2 and III.3, respectively.

### v2

@@ Abstract
Quantum entanglement in strongly correlated electron systems often leads to exotic elementary excitations. Quantum spin liquids provide a paradigmatic example, where the elementary excitations are described by fractional quasiparticles such as spinons. However, such fractional quasiparticles behave differently from electrons, making their experimental identification challenging. Here, we theoretically investigate the spin Seebeck effect, which is a thermoelectric response via a spin current, as an efficient probe of the fractional quasiparticles in quantum spin liquids, focusing on the Kitaev honeycomb model. By comprehensive studies using the real-time dynamics, the perturbation theory, and the linear spin-wave theory based on the tunnel spin-current theory, we find that the spin current is induced by thermal gradient in the Kitaev spin liquid, via the low-energy fractional Majorana excitations. This underscores the ability of Majorana fermions to carry spin current, despite lacking spin angular momentum. Furthermore, we find that the induced spin current changes its sign depending on the sign of the Kitaev interaction, indicating that the Majorana fermions contribute to the spin current with (up-)down-spin-like nature when the exchange coupling is (anti)ferromagnetic. Thus, in contrast to the negative spin current already found in a one-dimensional quantum spin liquid, our finding reveals that the spin Seebeck effect can exhibit either positive or negative signals, contingent upon the nature of fractional excitations in the quantum spin liquids. We also clarify contrasting field-angle dependence between the Kitaev spin liquid in the low-field limit and the high-field ferromagnetic state, which is useful for the experimental identification. Our finding suggests that the spin Seebeck effect could be used not only to detect fractional quasiparticles emerging in quantum spin liquids but also to generate and control them.
@@ I Introduction
Quantum entanglement is the key to the emergence of topological phases which have been of central interest in strongly correlated electron systems. The topological phases can be regarded as vacuums of fractionalized elementary excitations which behave differently from the original electrons. A representative example is the fractional quantum Hall states in two-dimensional electron systems under a magnetic field [1], where electrons and magnetic fluxes form composite elementary excitations with fractional charges obeying neither the Bose-Einstein nor Fermi-Dirac statistics, called non-Abelian anyons [2, 3, 4]. Numerous studies have been conducted to identify and manipulate the non-Abelian anyons [5], since they are expected to be useful for future quantum technologies, such as fault-tolerant quantum computing [6, 7].
@@
The Kitaev spin liquid (KSL) is a rare quantum spin liquid that has been rigorously shown to be the ground state of a realistic quantum spin model on a two-dimensional honeycomb lattice [24]. The exact solution allows one to scrutinize the fundamental properties of fractional excitations. In the KSL, spins are fractionalized into itinerant Majorana fermions and localized $Z_{2}$ fluxes at zero magnetic field, and these two form a composite obeying anyonic statistics under a magnetic field. Stimulated by the proposal for realization of the model [25], tremendous efforts have been devoted to exploring candidate materials [26, 27, 28]. In addition, close collaborations between experiment and theory have unveiled exotic properties of the KSL [29, 30, 26, 31, 32]. A striking finding is the half-integer quantized thermal Hall effect as a direct evidence of the fractional Majorana excitations [33, 34], while the experimental results are still controversial [35, 36, 37, 38]. It was pointed out that the sample quality matters [39]. Significant contributions from phonons were also discussed [40, 41]. Recently, there are several theoretical proposals for generating, detecting, and controlling the fractional excitations in the KSL, e.g., by using the scanning tunneling microscopy [42, 43, 44, 45, 46], interferometers [47, 48, 49], impurity effects [50, 51], local lattice distortions [52], and magnetic fields [53], whose experimental demonstrations are awaited.
@@
Figure 1: (a) Schematic of an experimental setup for measuring the spin Seebeck effect in Kitaev magnets. A Kitaev magnet is attached to heatbaths from two sides to generate a temperature gradient ($-\nabla T$). A metal with strong spin-orbit coupling is sandwiched at one end. Due to the inverse spin Hall effect, the spin current is converted into an electric field $E_{\rm ISHE}$ (light blue arrow) perpendicular to the thermal gradient in the metal, which is observable. The tunnel spin current $\tilde{J}_{S}$ (red arrow) and the magnetic field ${\bf B}$ parallel to $c$ (green arrow) are also depicted. We define the sign of $\tilde{J}_{S}$ to be positive when the tunnel spin current polarized antiparallel to ${\bf B}$ direction flows from the hot side to the cold side. Actual direction of the tunnel spin current depends on the sign of the Kitaev interaction; see Sec. III. (b) Kitaev model on a honeycomb lattice used in simulations. The black dots represent the spin-1/2 degrees of freedom, and the red, green, and blue lines represent the $x$, $y$, and $z$ bonds, respectively [see Eq. (1)]. The lattice with $X=4$ and $L=12$ is shown, where $X$ is the number of spins in the vertical direction and $L$ is that in the horizontal direction. We impose an open boundary condition in the horizontal direction and a periodic boundary condition in the vertical direction. The $a$, $b$, and $c$ axes and $S^{x}$, $S^{y}$, and $S^{z}$ axes are shown in the inset. $j_{0}$ denotes the site where the local dynamical spin susceptibility is calculated.
@@
The spin Seebeck effect (SSE) refers to a phenomenon in which a spin current is induced by a thermal gradient. This effect is widely used as a sensitive probe for elementary excitations in magnets [54, 55, 56, 57, 58, 59, 60, 61]. In fact, in ordinary ferromagnets and antiferromagnets, the SSE is induced by magnon excitations from the conventional magnetic ordered state. Recently, this probe has been extended to more nontrivial magnets, such as those with frustrated interactions and low dimensionality, whose elementary excitations may have both positive and negative spin angular momenta, or even become spinless. The representative example is a negative SSE observed in a quantum spin liquid in a quasi-one-dimensional antiferromagnet [62]. This peculiar behavior, in contrast to conventional ferromagnets, was ascribed to fractional spinon excitations forming the Tomonaga–Luttinger liquid. Besides, other systems with strong quantum fluctuations, such as a spin-nematic liquid [63], a spin-Peierls system [64], and a magnon Bose–Einstein condensation system [65], have also been reported to exhibit the peculiar SSE. The SSE was also theoretically predicted for the KSL [66]. However, this is caused by chiral edge modes in the KSL, and no study of bulk SSE for quantum spin liquids in more than one dimension has been reported thus far to the best of our knowledge. Recently, unconventional spin transport properties to thermal conductivity mediated by Majorana excitations in the KSL have been reported experimentally [67] and theoretically [68, 69, 70]. Nonetheless, it remains unclear how the SSE, which involve both spin and heat transport, is manifested in the KSL and whether it is useful to identify, generate, and control the fractional excitations.
@@
In this paper, we theoretically analyze the spin current induced by a thermal gradient via the SSE in the KSL. The possible experimental setup is shown in Fig. 1(a); we consider an interface between a Kitaev magnet and a paramagnetic metal, and calculate the tunnel spin current through the interface by the local dynamical spin susceptibility in an applied magnetic field, based on the tunnel spin-current theory [71, 72]. Accurately calculating the dynamics of quantum many-body problems, particularly in more than one spatial dimension, poses a significant challenge. In addressing this issue, we opt for a state-of-the-art numerical technique, real-time dynamics simulations based on the time-dependent variational principle (TDVP) [73] with matrix product states (MPSs) [74, 75]. We also complementarily conduct the analysis based on the perturbation theory with respect to the magnetic field for the KSL, as well as the linear spin-wave theory for the field-induced ferromagnetic (FFM) state. As a result, we find that the spin current is induced by a thermal gradient in the KSL mediated by the low-energy fractional Majorana quasiparticles that have no spin angular momentum. This is qualitatively different from the fact that fractional spinons in forming a one-dimensional Tomonaga–Luttinger liquid induce a negative SSE since the spinon has spin-$1/2$. It is remarkable that the quasiparticles without angular momentum can contribute to the tunnel spin-current transport. Furthermore, we find that the sign of the tunnel spin current depends on the sign of the Kitaev interaction; namely, the ferromagnetic (FM) and AFM KSLs lead to a positive and negative SSE, respectively. This is in stark contrast to the behavior in conventional ferromagnetic states where the magnon excitations lead to a positive SSE. In addition, we find that the KSLs show distinctive field angle dependence of the spin current that is different from those of the FFM states. These peculiar behaviors, which are ascribed to the spin current mediated by fractional Majorana excitations, are revealed for the first time to our knowledge by using the state-of-the-art numerical technique. In turn, our findings suggest the possibility of generating and controlling the Majorana excitations by spin injection.
@@
The structure of this paper is as follows. In Sec. II, we describe the model and method used in this study. After introducing the Kitaev model under an external magnetic field, we describe the tunnel spin-current theory and numerical methods of real-time dynamics simulations to compute the SSE. We also mention the choice of model parameters and the details of calculation conditions. In Sec. III, we present the results of the real-time dynamics simulations, the perturbation analysis, and the linear spin-wave theory. In Sec. IV, we discuss the contrasting behaviors of the tunnel spin current depending on the sign of the Kitaev interaction, and the strength and direction of the magnetic field, which are important for experimental confirmation, and the experimental implication and observability of the SSE in the KSL. Finally, Sec. V is devoted to summary.
@@ II Model and method
We investigate the SSE in the Kitaev model in Eq. (1) based on the tunnel spin-current theory [71, 72, 64, 76]. The experimental setup is shown in Fig. 1(a); the SSE is measured by the inverse spin Hall effect [77, 78, 79] in a metal with strong spin-orbit coupling, such as Pt, attached to a Kitaev magnet. In the tunnel spin-current theory, spin injection is assumed to occur locally through the interface between the metal and the magnet, and calculated under the following conditions: (i) The interaction between the metal and the magnet is small enough to be dealt with as a perturbation and conserves angular momenta, (ii) the polarization of the tunnel spin current is parallel to the direction of the magnetic field, (iii) the temperature difference between the metal and the magnet is small, and (iv) the metal is nonmagnetic with featureless density of states in the low-energy region. It is worth noting that although the total magnetization is not conserved in the Kitaev model, these assumptions allow us to evaluate the tunnel spin current transferred to the metal that is detected in experiments.
On these assumptions, the tunnel spin current induced by a thermal gradient is evaluated by the perturbation theory with respect to the interface interaction using the nonequilibrium Green function method [80, 81, 82]. The tunnel spin current per one interface site is given by
@@
up to a constant coefficient proportional to the square of the interface interaction and the temperature difference between the metal and magnet. In this paper, the sign of the tunnel spin current is defined such that $\tilde{J}_{S}>0$ when the tunnel spin current polarized antiparallel to the magnetic field direction flows in the direction of the red arrow in Fig. 1(a). The local dynamical spin susceptibility of the magnet, $\chi^{-+}_{\rm loc}(\omega)$, is defined as
@@
with the damping factor $\delta$; $S^{\pm}_{j_{0}}$ is defined according to the direction of the magnetic field, namely,
@@
where $\tilde{x}$ and $\tilde{y}$ represent the directions perpendicular to the magnetic field along $\tilde{z}$ (the $\tilde{x}\tilde{y}\tilde{z}$ coordinate forms a right-handed system) 111 We define $j_{0}$ on only one of the two sublattices as shown in Fig. 1(b) and neglect the sublattice dependence because trial simulation results do not show recognizable difference in $\chi_{\rm loc}^{-+}(\omega)$ when $j_{0}$ is defined on the other sublattice. . In Eq. (2), the integral kernel $k(x=\omega/T)$ is given by
@@
In this study, we compute the tunnel spin current in an applied magnetic field ${\bf B}$ along the $\tilde{z}(=z$, $a$, $b$, or $c)$ axis [see Fig. 1(b)], and mainly discuss the cases of ${\bf B}\parallel z$ and ${\bf B}\parallel c$; the former case gives the Zeeman coupling as $-BS_{j}^{z}$, while the latter gives $-B(S_{j}^{x}+S_{j}^{y}+S_{j}^{z})/\sqrt{3}$ in the second term in Eq. (1). In the calculations of $X_{+}(\omega)$, we choose $S_{j_{0}}^{\pm}$ as $S_{j_{0}}^{\pm}=S_{j_{0}}^{x}\pm iS_{j_{0}}^{y}$ for ${\bf B}\parallel z$, and $S_{j_{0}}^{\pm}=S_{j_{0}}^{a}\pm iS_{j_{0}}^{b}$ for ${\bf B}\parallel c$ (cyclic permutations for ${\bf B}\parallel a$ and ${\bf B}\parallel b$). We investigate the low-field KSL as well as the high-field FFM state for comparison. The phase transition from the KSL to the FFM state was studied by using the exact diagonalization, the density matrix renormalization group (DMRG) [88, 89], and Majorana mean-field approximation [84, 85, 86, 87]. The phase diagram is different between the FM and AFM cases not only quantitatively but also qualitatively, as shown in Fig. 3. In the FM case, the KSL turns into the FFM state at $B\approx 0.03$ for both ${\bf B}\parallel c$ [84, 85, 87] and ${\bf B}\parallel z$ [86]; there is a single phase transition. Meanwhile, in the AFM case, the values of the critical field become one order of magnitude larger, and moreover, the system exhibits another phase transition in addition to that to the FFM state. For ${\bf B}\parallel c$, the KSL is stable up to $B\approx 0.4$, and a different spin liquid (SL) takes place before entering into the FFM state at $B\approx 0.6$ [85, 87]. For ${\bf B}\parallel z$, the KSL is stable up to $B\approx 0.4$, and the intermediate SL is predicted to be stable up to $B\approx 0.5$ [86]. With reference to these studies, we choose the values of $B$ for the KSL (FMM) state, $B_{\rm KSL}$ ($B_{\rm FFM}$), well below (above) the critical fields: $B_{\rm KSL}=0.01$ and $B_{\rm FFM}=0.5$ for the FM case, and $B_{\rm KSL}=0.1$ and $B_{\rm FFM}=1.5$ for the AFM case, as indicated by the red arrows in Fig. 3.
@@
In the following calculations, we compute $\chi^{-+}_{\rm loc}(\omega)$ in Eq. (3) at zero temperature for simplicity. This leaves the temperature dependence of the tunnel spin current only in the kernel $k(\omega/T)$. This assumption should be justified at sufficiently low temperature where the $Z_{2}$ flux excitations are scarce, typically below the crossover temperature $T\approx 0.016$ at zero field [90]. To compute $\chi^{-+}_{\rm loc}(\omega)$ at $T=0$, we perform a real-time dynamics simulations based on TDVP [91]. The initial states of the simulations are obtained by multiplying $S^{\pm}_{j_{0}}$ to the MPS ground states computed by DMRG with the bond dimension $400$. We consider a cylinder-shaped lattice structure with the periodic (open) boundary condition in the vertical (horizontal) direction, as shown in Fig. 1(b), of the system size $X\times L$ with $(X,L)=(4,12)$, $(4,16)$, $(6,6)$, $(6,8)$, and $(6,10)$, where $X$ and $L$ are the number of sites in the vertical and horizontal directions, respectively, and employ the MPS representation wrapping around the cylinder in a snake form. We note that the ground state is not degenerate for all cases at the chosen magnetic fields. In Eq. (3), we choose $j_{0}$ at the site close to the center of each cylinder, as shown in Fig. 1(b), and take $\delta=0.01$. We set one time step of real-time dynamics as $\Delta t=0.2$ and perform $N_{t}=10^{4}$ time evolutions: Simulations are performed up to a maximum time $t_{\rm max}=2000$. When performing the Fourier transform [Eq. (3)], we interpolate the data by a cubic spline. The calculations are performed using a software library iTensor [92]. In addition, to understand the simulation results, we analyze the dynamical spin susceptibility based on the perturbation theory with respect to the magnetic field and the linear spin-wave theory, as described in Secs. III.2 and III.3, respectively.